A thin, horizontal copper rod is 1.20 m long and has a mass of 44.5g. What is the minimum current in the rod that can cause it to float in a horizontal magnetic field of 2.45 T?
Answer #1

BIL = mg

=> 2.45 * I * 1.2= 44.5 * 10^(-3) * 9.81

=> I =0.1484 A
Answer #2

a similar question is solved below, but with different values. hope this helps you.

A thin horizontal copper rod is 1.0 m long and has a mass of 65.0 g. What is the minimum current in the rod that can cause it to float in a horizontal magnetic field of 2.0 T?


To make the rod float, the weight of rodmust be balanced by the magnetic force on it.
m *g = B * I * L * sin?
If the current is to be minimum, sin ?must be maximum i.e. 1.
=> 65 * 10-3* 9.8 = 2.0 * I * 1.0

I = 6.37 *10-1 /2.0 = 0.3185 A
Answer #3
To make the rod float, the weight of rodmust be balanced by the magnetic force on it.
m *g = B * I * L * sin?
If the current is to be minimum, sin ?must be maximum i.e. 1.
=> 65 * 10-3* 9.8 = 2.0 * I * 1.0